OpenStudy (anonymous):

simplify: (-7)+6+[-(2-3)]

OpenStudy (luigi0210):

Do what's in the brackets first, then work from left to right.

OpenStudy (anonymous):

pemdas parenthesis,exponets, multiply,divide, add subtract. you do whats in brackets first left to right. -2--3=-2+3=-1 -7+6=-1 -1=-1 ALL REAL NUMBERS cursive-ish R